# Dirac Equation Solution with Generalized tanh-Shaped Hyperbolic Potential: Application to Charmonium and Bottomonium Mass Spectra

### v1

@@ Abstract
In this work, we advance the understanding of the fundamental interactions by using a generalized tanh-shaped hyperbolic potential to investigate bound state solutions of the Dirac equation. Using the Nikiforov-Uvarov method, we derive energy eigenvalues and radial wave functions in various quantum states, with the results expressed in terms of hypergeometric polynomials. The results of our analysis indicate that the energy eigenvalues are strongly correlated with the potential parameters and covered several fundamental potential models used to expose a non-trivial behaviour that occurred within quantum systems. Further, using this potential to model mass spectra of charmonium and bottomonium, we show that results for the calculated quark mass spectra align closely with experimentally observed values, reflecting the veracity and accuracy of the approach. This close correspondence between calculated and experimental spectra highlights this potential model’s efficacy in describing heavy quarkonium systems. The insights gained from this study reinforce the applicability of this potential model in quantum mechanics and particle physics and lay a foundation for its extension to other quantum systems, suggesting a feasible avenue for future research.
@@ I Introduction
[…]
[findent=2pt]The Dirac equation (DE) stands as a fundamental cornerstone within the intricate framework of quantum mechanics.Dirac and Bohr (1927); Dirac (1963) This equation is the deep one in to the conduct of fermions at high-relativistic velocities.Dirac and Bohr (1927); Dirac (1963); Feshbach and Villars (1958); Dirac (1949); Chirgwin and Flint (1945); Meng et al. (1998); Campos and Fabbri (2022); Bagrov et al. (1975); Campos and Cabrera (2020); Barnett (2017); Campos et al. (2017); Bialynicki-Birula (2004); Kaminer et al. (2015); Hammad et al. (2024); Bialynicki-Birula (2021) This wave equation, bridging between quantum mechanics and special relativity, has revolutionized particle physics as well as retained fertile grounds for significant exploration till date.Giuliani et al. (2019); Li et al. (2016); Ding et al. (2021); Agarwal (2019)
The DE is at the core of the concepts of spin and pseudo-spin symmetries, which are embedded within the Dirac Hamiltonian.Alberto et al. (2001); Ginocchio (1997, 1999, 2004, 2005); Liang et al. (2015); Zhou et al. (2003); Page et al. (2001); Arima et al. (1969); Hecht and Adler (1969) They explain a variety of physical phenomena associated with the simplicial background and are definitive tools in the interpretation of nuclear structure and reactions.Ginocchio (1997, 1999, 2004, 2005); Liang et al. (2015); Zhou et al. (2003) Spin symmetry, manifesting as a spin doublet, leads to degeneracy between states characterized by quantum numbers $(n,l,j=l\pm s)$Ginocchio (2004, 2005); Liang et al. (2015); Zhou et al. (2003). This symmetry is instrumental in elucidating the spectrum of antinucleons in a nucleus and the subtle spin-orbit splitting observed in hadronsPage et al. (2001). In contrast, pseudo-spin symmetry, another form of degeneracy, arises between states with quantum numbers $(n,l,j=l+\frac{1}{2})$ and $(n-1,l+2,j=l+\frac{3}{2})$Arima et al. (1969); Hecht and Adler (1969). This symmetry has profound implications for understanding nuclear deformation de Voigt et al. (1983); Dudek et al. (1987), identical bands,Nazarewicz et al. (1990); Zeng et al. (1991); Byrski et al. (1990) magnetic moments Troltenier et al. (1994); Stuchbery (2002), magic number shifts Nagai et al. (1981); Gaudefroy et al. (2006); Bastin et al. (2007) and effective shell-model structures in nuclear physics.Troltenier et al. (1995)
Pseudo-spin symmetry was initially explored within the non-relativistic framework and later extended into relativistic mean field theory.Page et al. (2001); Arima et al. (1969) Significant strides have been made in this domain, with comprehensive investigations revealing the intricate details of spin and pseudo-spin symmetries.Ginocchio (2005); Liang et al. (2015); Ahmadov et al. (2022) Significant advancements have been made in this area, particularly through the analytical solutions of the DE with physical potentials, which are crucial for uncovering the full spectrum of a quantum system’s properties de Castro and Alberto (2012); Xu et al. (2015); Fang et al. (2017); Hartmann and Portnoi (2014). In this context, our focus is on solving the DE for the generalized tanh-shaped hyperbolic potential (GTHP) Ahmadov et al. (2021); Badalov and Badalov (2023), described as:
@@
where $V_{1}$, $V_{2}$, and $V_{3}$ represent potential well depths, and $\alpha$ characterizes the interaction potential’s properties. The GTHP exhibits a minimum value $V_{\text{min}}=V(r_{\text{e}})=V_{1}-\frac{{V_{2}}^{2}}{4V_{3}}$ at $r_{\text{e}}=\frac{1}{\alpha}\tanh^{-1}\left(-\frac{V_{2}}{2V_{3}}\right)$, subject to $\left|V_{2}\right|<2V_{3}$ and $V_{3}>0$.Ahmadov et al. (2021); Badalov and Badalov (2023) This potential encompasses a wide range of physical scenarios, including various standard and generalized potential models, including the standard and generalized Woods-Saxon,Alberto et al. (2001); Woods and Saxon (1954); Badalov et al. (2019) Rosen-Morse,Rosen and Morse (1932) Manning-Rosen type,Manning and Rosen (1933) standard and generalized Morse (improved Rosen-Morse),Morse (1929); Jia et al. (2012) Schiöberg,Schiöberg (1986); Wang et al. (2012) four-parametric exponential type,Chun-Sheng et al. (2001); Ikot et al. (2016) Williams-Poulios,Williams and Poulios (1993); Peña et al. (2017) and the sum of the linear and harmonic oscillator potentials in some special casesAhmadov et al. (2021); Badalov and Badalov (2023). Notably, considering $V_{1}\neq 0$ in GTHP facilitates the modeling of these particular casesAhmadov et al. (2021); Badalov and Badalov (2023). These potentials are particularly relevant in modeling complex quantum systems, where the GTHP’s flexibility allows it to capture a broad spectrum of interactions.
@@
In this work, we explore the bound-state solutions of the DE within the framework of the GTHP. Employing the Nikiforov-Uvarov (NU) method, we derive exact expressions for the energy eigenvalues and the corresponding radial wave functions articulated through hypergeometric polynomials for various quantum states. Our analysis reveals the sensitivity of these eigenvalues to variations in the potential parameters, providing deep insights into the quantum dynamics encapsulated by the GTHP model. Furthermore, we extend our theoretical framework to studying heavy quark systems, specifically targeting the charmonium and bottomonium mass spectra. Heavy quarkonia, consisting of a heavy quark and antiquark (such as $b\bar{b}$ and $c\bar{c}$), serve as a critical testing ground for quantum chromodynamics and various potential models.Patel and Vinodkumar (2009); Brambilla et al. (2005); Purohit et al. (2022); Soni et al. (2018); Bukor and Tekel (2023); Li and Chao (2009); Ni et al. (2022) These systems are particularly significant owing to their rich spectroscopy, with many states lying below the threshold of open charm or bottom production.Brambilla et al. (2005); Purohit et al. (2022); Soni et al. (2018); Bukor and Tekel (2023); Li and Chao (2009); Ni et al. (2022); Eichten and Quigg (1995); Kawanai and Sasaki (2014); Godfrey and Isgur (1985); Quigg and Rosner (1979); Brambilla et al. (2011); Eichten et al. (1978); Sang et al. (2023); Particle Data Group et al. (2022); Godfrey and Moats (2015); Mishra et al. (2024) Employing the GTHP model to these systems, one is able to extract bound-state masses with quite good agreement with experimental results leading to important knowledge on heavy quark interaction dynamics and properties of quark-antiquark potentials on hadronic scales.
@@ II Solutions to the Radial Dirac Equation
The DE serves as a fundamental framework for describing the quantum behavior of fermions in the presence of scalar and vector potentials.Greiner (2001) For a particle interacting with an attractive scalar potential $S(\vec{r})$ and a repulsive vector potential $V(\vec{r})$, the DE is given by:
@@
where $E$ is the relativistic total energy, $\hat{\vec{p}}=-i\hbar\vec{\nabla}$ is the momentum operator, and $\vec{\alpha}$ and $\beta$ are the standard Dirac matrices.Greiner (2001) To study the system under the influence of the GTHP, we recast the DE into its radial form, exploiting spherical symmetry. In this context, both potentials $S(r)$ and $V(r)$ are functions of the radial coordinate $r$, which allows us to decompose the Dirac spinor into radial and angular parts. Consequently, the DE simplifies to coupled radial differential equations for the upper and lower components of the wave function. This formalism is particularly efficient for investigating energy eigenvalues and wave functions in spherically symmetric potentials.
For spin and pseudo-spin symmetry cases, the corresponding energy eigenvalues $E_{nk}$ and normalized radial wave functions are derived. The details of these derivations are discussed in Ref. [SM, ], where specific symmetry conditions simplify the DE, such as $\Delta(r)=C_{s}$ for spin symmetry and $\Sigma(r)=C_{ps}$ for pseudo-spin symmetry.
@@
$\quad\left.+\frac{\frac{\left(Mc^{2}+E_{nk}-C_{s}\right)V_{2}}{2{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k+1)}{2{\alpha}^{2}r^{2}_{e}}A_{1}}{\sqrt{\frac{1}{4}+\frac{\left(Mc^{2}+E_{nk}-C_{s}\right)V_{3}}{{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k+1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}}\right)^{2}$
The information about all parameters in Eq. (S32) and the full derivation process to obtain Eq. (S32) are provided in Section S1 of Ref. [SM, ].
@@
$\quad\left.+\frac{-\frac{(Mc^{2}-E_{nk}+C_{ps})V_{2}}{2{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k-1)}{2{\alpha}^{2}r^{2}_{e}}A_{1}}{\sqrt{\frac{1}{4}-\frac{(Mc^{2}-E_{nk}+C_{ps})V_{3}}{{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k-1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}}\right)^{2}$
The information about all parameters in Eq. (S49) and the full derivation process to obtain Eq. (S49) are provided in Section S1 of Ref. [SM, ].
@@ III.1 Investigation of Spin Symmetry Case in Specific Scenarios
The GTHP is a highly versatile framework that exhibits a remarkable sensitivity to its parameters, making it an effective model to describe several prominent potentials encountered in fundamental quantum systems. The parametric flexibility of GTHP allows it to emulate various well-known potentials, such as the generalized Woods-Saxon,Woods and Saxon (1954); Badalov et al. (2019) Rosen-Morse,Rosen and Morse (1932) and Manning-Rosen potentials,Manning and Rosen (1933) among others. This adaptability stems from the intricate control of key parameters, such as $V_{1}$, $V_{2}$, $V_{3}$, and $\alpha$, which dictate the potential’s shape, depth, and range.
It is well-known that spin symmetry occurs in the DE when the difference between the vector and scalar potentials is constant, $V(r)-S(r)=\text{constant}$. In this regime, the spin-orbit coupling becomes negligible, leading to near-degenerate energy levels and simplifying the relativistic energy spectrum.Ginocchio (1997, 1999) The ability to smoothly transition between different potential profiles by tuning these parameters makes GTHP an ideal model for exploring the underlying physics of systems with spin symmetry. Each set of parameters generates a distinct potential landscape, which strongly influences the system’s energy levels, the degree of degeneracy, and the manifestation of spin symmetry. This symmetry is significant in high-energy physics and nuclear physics, particularly in understanding mesonic and baryonic spectra, as well as nucleon-nucleon interactions. The sensitivity of GTHP model is also crucial for tailoring specific quantum systems and analyzing the behavior of relativistic particles in various physical environments as well as allowing for the emulation of short-range and long-range interactions across various fields of fundamental physics. We now examine the energy eigenvalue derived from Eq. (S32) for different potentials modeled under the GTHP framework.
@@
$C_{2}$ $=\frac{4A_{2}}{{\left(1+x_{e}\right)}^{2}}.$
@@
ii) Considering the case when $W=0$ and $x_{e}=0$ for the energy levels equation of the standard Woods-Saxon potential Feizi et al. (2012), we derive:
@@
iii) For the energy spectrum equation of the Rosen-Morse potential with parameters $\ V_{3}=-V_{1}=C$ and $V_{2}=B$, the formulation is as follows:
@@
$\left.-\frac{\frac{(Mc^{2}+E_{nk}-C_{s})(W_{1}-W_{3})}{4{\alpha}^{2}{\hslash}^{2}c^{2}}-\frac{k(k+1)}{2{\alpha}^{2}r^{2}_{e}}A_{1}}{\sqrt{\frac{1}{4}+\frac{(Mc^{2}+E_{nk}-C_{s})(W_{1}-W_{2}+W_{3})}{4{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k+1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}}\right)^{2}$
@@
×$\times\left(n+\frac{1}{2}\right)-{\alpha}^{2}{\hslash}^{2}c^{2}\left[\left(1+\frac{3V^{2}_{2}}{4V^{2}_{3}}\right){\left(n+\frac{1}{2}\right)}^{2}+\frac{1}{4}\left(1-\frac{V^{2}_{2}}{4V^{2}_{3}}\right)\right]+o\left({\alpha}^{2}\right)=0$
@@
x) Setting $C_{s}=0$ allows us to simplify the potential for the energy level equation of a Klein-Gordon particle, resulting in Ref. [Badalov and Badalov, 2023]:
@@
$\left.+\frac{\frac{(Mc^{2}+E_{nk})V_{2}}{2{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{l(l+1)}{2{\alpha}^{2}r^{2}_{e}}A_{1}}{\sqrt{\frac{1}{4}+\frac{(Mc^{2}+E_{nk})V_{3}}{{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{l(l+1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}}\right)^{2}=0$
@@
$\left.+\frac{\frac{\mu V_{2}}{{\alpha}^{2}{\hslash}^{2}}+\frac{l(l+1)}{2{\alpha}^{2}r^{2}_{e}}A_{1}}{\sqrt{\frac{1}{4}+\frac{2\mu V_{3}}{{\alpha}^{2}{\hslash}^{2}}+\frac{l(l+1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}}\right]^{2}$
This formulation aligns with the findings presented in Ref. [Ahmadov et al., 2021], valid for $n=0,1,2,\ldots,n_{max}$.

### v4

@@ Abstract
In this study, we present an analytical solution of the Dirac equation in a generalized tanh-shape hyperbolic potential, which allows us to unify various well-known quantum potentials under a single theoretical framework. This versatile potential model is used to compute the mass spectra for charmonium and bottomonium, with excellent agreement with experimental measurements and does better than some potential models in predicting the several orbital states. Our results not only validate the GTHP as a powerful tool for describing heavy quarkonium systems but also suggest its broader applicability in exploring quantum systems where similar potentials are effective. This work is a stepping stone for new research into fermionic systems with complex interactions, by jointly providing insights into foundational aspects of quantum mechanics as well as applications in particle physics.
@@ I Introduction
The Dirac equation (DE) is a cornerstone of quantum mechanics and is responsible for revolutionizing theoretical physics, giving rise to what is in essence the first relativistic quantum field theory, leading to the prediction of novel phenomena such as antimatter [1]. The equation has led to advances in many areas in physics, from quantum electrodynamics [2] to the way particles behave in the relativistic limit [3, 4]. Responsible in its particular nuclear properties studies [5], which later extended to be used in potent modern effort in a superheavy element [6], complex quantum systems [7], chiral symmetric quantum chromodynamics [8] and antiparticles [9]. In addition, DE continues to be relevant for current research and opens the possibility to study electron vortices [10], coherent control [11] and new theoretical approaches [12, 13, 14, 15, 16].
The DE is at the core of the concepts of spin and pseudo-spin symmetries, which are embedded within the Dirac Hamiltonian [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. They explain a variety of physical phenomena associated with the simplicial background and are definitive tools in the interpretation of nuclear structure and reactions [18, 19, 20, 21, 22, 23]. Spin symmetry, manifesting as a spin doublet, leads to degeneracy between states characterized by quantum numbers $(n,l,j=l\pm s)$[20, 21, 22, 23]. This symmetry is instrumental in elucidating the spectrum of antinucleons in a nucleus and the subtle spin-orbit splitting observed in hadrons [24]. In contrast, pseudo-spin symmetry, another form of degeneracy, arises between states with quantum numbers $(n,l,j=l+\frac{1}{2})$ and $(n-1,l+2,j=l+\frac{3}{2})$[25, 26]. This symmetry has profound implications for understanding nuclear deformation [27, 28], identical bands [29, 30, 31], magnetic moments [32, 33], magic number shifts [34, 35, 36] and effective shell-model structures in nuclear physics [37].
Pseudo-spin symmetry was initially explored within the non-relativistic framework and later extended into relativistic mean field theory [24, 25]. Significant strides have been made in this domain, with comprehensive investigations revealing the intricate details of spin and pseudo-spin symmetries [21, 22, 38]. Significant advancements have been made in this area, particularly through the analytical solutions of the DE with physical potentials, which are crucial for uncovering the full spectrum of a quantum system’s properties [39, 40, 41, 42]. In this context, our focus is on solving the DE for the generalized tanh-shaped hyperbolic potential (GTHP) [43, 44], described as:
@@
where $V_{1}$, $V_{2}$, and $V_{3}$ represent potential well depths, and $\alpha$ characterizes the interaction potential’s properties. The GTHP exhibits a minimum value $V_{\text{min}}=V(r_{\text{e}})=V_{1}-\frac{{V_{2}}^{2}}{4V_{3}}$ at $r_{\text{e}}=\frac{1}{\alpha}\tanh^{-1}\left(-\frac{V_{2}}{2V_{3}}\right)$, subject to $\left|V_{2}\right|<2V_{3}$ and $V_{3}>0$[43, 44]. This potential encompasses a wide range of physical scenarios, including various standard and generalized potential models, including the standard and generalized Woods-Saxon [17, 45, 46], Rosen-Morse [47], Manning-Rosen type [48], standard and generalized Morse (improved Rosen-Morse) [49, 50], Schiöberg [51, 52], four-parametric exponential type [53, 54], Williams-Poulios [55, 56], and the sum of the linear and harmonic oscillator potentials in some special cases [43, 44]. Notably, considering $V_{1}\neq 0$ in GTHP facilitates the modeling of these particular cases[43, 44]. These potentials are particularly relevant in modeling complex quantum systems, where the GTHP’s flexibility allows it to capture a broad spectrum of interactions.
@@
While previous studies have significantly advanced our understanding of so-called spin symmetries in relativistic and non-relativistic fields and clarified the critical spin dynamics, there is still a significant gap in applying these insights to complex potentials that more accurately reflect real quantum phenomena. With this study, we aim to fill this gap by advancing our understanding of the hadronic physics and GTHP. This work investigates the bound-state solutions of the DE from the perspective of the GTHP. Employing the Nikiforov-Uvarov (NU) method, reducing the second-order differential equation to a hypergeometric form, we derive exact expressions for the energy eigenvalues and the corresponding radial wave functions articulated through hypergeometric polynomials for various quantum states. Our analysis reveals the sensitivity of these eigenvalues to variations in the potential parameters, providing deep insights into the quantum dynamics encapsulated by the GTHP model. Furthermore, we extend our theoretical framework to studying heavy quark systems, specifically targeting the charmonium and bottomonium mass spectra. Heavy quarkonia, consisting of a heavy quark and antiquark (such as $b\bar{b}$ and $c\bar{c}$), serve as a critical testing ground for quantum chromodynamics (QCD) and various potential models [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67]. These systems are particularly significant owing to their rich spectroscopy, with many states lying below the threshold of open charm or bottom production [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 16]. Employing the GTHP model to these systems, one is able to extract bound-state masses with quite good agreement with experimental results leading to important knowledge on heavy quark interaction dynamics and properties of quark-antiquark potentials on hadronic scales.
@@ II Solutions to the Radial Dirac Equation
The DE serves as a fundamental framework for describing the quantum behavior of fermions in the presence of scalar and vector potentials [77]. For a particle interacting with an attractive scalar potential $S(\vec{r})$ and a repulsive vector potential $V(\vec{r})$, the DE is given by:
@@
where $E$ is the relativistic total energy, $\hat{\vec{p}}=-i\hbar\vec{\nabla}$ is the momentum operator, and $\vec{\alpha}$ and $\beta$ are the standard Dirac matrices [77]. To study the system under the influence of the GTHP, we recast the DE into its radial form, exploiting spherical symmetry. In this context, both potentials $S(r)$ and $V(r)$ are functions of the radial coordinate $r$, which allows us to decompose the Dirac spinor into radial and angular parts. Consequently, the DE simplifies to coupled radial second-order linear differential equations for the upper and lower components of the wave function. This formalism is particularly efficient for investigating energy eigenvalues and wave functions in spherically symmetric potentials. Then, we applied NU method [78] to solve this radial second-order linear differential equations, see the derivation which is discussed detailed in Ref. [79]. The transition to a radial form is crucial as it highlights the symmetries inherent in DE, especially when analyzing the effects of scalar and vector potentials that vary only with radial distance. By effectively isolating the radial components, this approach facilitates a focused study of the impact of potential variations under global symmetries on fermionic behavior.
For spin and pseudo-spin symmetry cases, the corresponding energy eigenvalues $E_{nk}$ and normalized radial wave functions are derived. The details of these derivations are discussed in Ref. [79], where specific symmetry conditions simplify the DE, such as $\Delta(r)=C_{s}$ for spin symmetry and $\Sigma(r)=C_{ps}$ for pseudo-spin symmetry.
@@
$\quad\left.+\frac{\frac{\left(Mc^{2}+E_{nk}-C_{s}\right)V_{2}}{2{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k+1)}{2{\alpha}^{2}r^{2}_{e}}A_{1}}{\sqrt{\frac{1}{4}+\frac{\left(Mc^{2}+E_{nk}-C_{s}\right)V_{3}}{{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k+1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}}\right)^{2}.$
The information about all parameters in Eq. (S32) and the full derivation process to obtain Eq. (S32) are provided in Section S1 of Ref. [79].
@@
$\quad\left.+\frac{-\frac{(Mc^{2}-E_{nk}+C_{ps})V_{2}}{2{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k-1)}{2{\alpha}^{2}r^{2}_{e}}A_{1}}{\sqrt{\frac{1}{4}-\frac{(Mc^{2}-E_{nk}+C_{ps})V_{3}}{{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k-1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}}\right)^{2}.$
The information about all parameters in Eq. (S49) and the full derivation process to obtain Eq. (S49) are provided in Section S1 of Ref. [79].
While the current study is based on the NU method, we would like to note that recent developments in gauge theory and quantum integrability offer alternative frameworks for exact spectral analysis. For example, gauge theory approaches using the ordinary differential equation/integrable model correspondences have been applied to solve eigenvalue problems in N=2{N}=2 supersymmetric systems [80, 81]. Similarly, quantum integrability techniques based on functional Bethe ansatz equations have provided exact quantization conditions in contexts such as black hole quasi-normal mode spectra [82, 83]. These advances, including new exact solutions in supersymmetric gauge theory and gravity contexts, highlight the power of integrability-based methods and point to promising directions for extending the current model.
@@ III.1 Investigation of Spin Symmetry Case in Specific Scenarios
The GTHP is a highly versatile framework that exhibits a remarkable sensitivity to its parameters, making it an effective model to describe several prominent potentials encountered in fundamental quantum systems. The parametric flexibility of GTHP allows it to emulate various well-known potentials, such as the generalized Woods-Saxon [45, 46], Rosen-Morse [47], and Manning-Rosen potentials [48], among others. This adaptability stems from the intricate control of key parameters, such as $V_{1}$, $V_{2}$, $V_{3}$, and $\alpha$, which dictate the potential’s shape, depth, and range.
It is well-known that spin symmetry occurs in the DE when the difference between the vector and scalar potentials is constant, $V(r)-S(r)=\text{constant}$. In this regime, the spin-orbit coupling becomes negligible, leading to near-degenerate energy levels and simplifying the relativistic energy spectrum [18, 19]. The ability to smoothly transition between different potential profiles by tuning these parameters makes GTHP an ideal model for exploring the underlying physics of systems with spin symmetry. Each set of parameters generates a distinct potential landscape, which strongly influences the system’s energy levels, the degree of degeneracy, and the manifestation of spin symmetry. This symmetry is significant in high-energy physics and nuclear physics, particularly in understanding mesonic and baryonic spectra, as well as nucleon-nucleon interactions. The sensitivity of GTHP model is also crucial for tailoring specific quantum systems and analyzing the behavior of relativistic particles in various physical environments as well as allowing for the emulation of short-range and long-range interactions across various fields of fundamental physics. We now examine the energy eigenvalue derived from Eq. (S32) for different potentials modeled under the GTHP framework.
@@
$C_{2}$ $=\frac{4A_{2}}{{\left(1+x_{e}\right)}^{2}},$
@@
ii) Considering the case when $W=0$ and $x_{e}=0$ for the energy levels equation of the standard Woods-Saxon potential [84], we derive:
@@
iii) For the energy spectrum equation of the Rosen-Morse potential with parameters $V_{3}=-V_{1}=C\$and $V_{2}=B\$, the formulation is as follows:
@@
$\left.-\frac{\frac{(Mc^{2}+E_{nk}-C_{s})(W_{1}-W_{3})}{4{\alpha}^{2}{\hslash}^{2}c^{2}}-\frac{k(k+1)}{2{\alpha}^{2}r^{2}_{e}}A_{1}}{\sqrt{\frac{1}{4}+\frac{(Mc^{2}+E_{nk}-C_{s})(W_{1}-W_{2}+W_{3})}{4{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{k(k+1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}}\right)^{2}.$
@@
$\times\left(n+\frac{1}{2}\right)-{\alpha}^{2}{\hslash}^{2}c^{2}\left[\left(1+\frac{3V^{2}_{2}}{4V^{2}_{3}}\right){\left(n+\frac{1}{2}\right)}^{2}+\frac{1}{4}\left(1-\frac{V^{2}_{2}}{4V^{2}_{3}}\right)\right]+o\left({\alpha}^{2}\right)=0.$
@@
x) Setting $C_{s}=0$ allows us to simplify the potential for the energy level equation of a Klein-Gordon particle, resulting in Ref. [44]:
@@
$\left.+\frac{\frac{(Mc^{2}+E_{nk})V_{2}}{2{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{l(l+1)}{2{\alpha}^{2}r^{2}_{e}}A_{1}}{\sqrt{\frac{1}{4}+\frac{(Mc^{2}+E_{nk})V_{3}}{{\alpha}^{2}{\hslash}^{2}c^{2}}+\frac{l(l+1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}}\right)^{2}=0.$
@@
$\left.+\frac{\frac{\mu V_{2}}{{\alpha}^{2}{\hslash}^{2}}+\frac{l(l+1)}{2{\alpha}^{2}r^{2}_{e}}A_{1}}{\sqrt{\frac{1}{4}+\frac{2\mu V_{3}}{{\alpha}^{2}{\hslash}^{2}}+\frac{l(l+1)}{{\alpha}^{2}r^{2}_{e}}A_{2}}-n-\frac{1}{2}}\right]^{2}.$
This formulation aligns with the findings presented in Ref. [43], valid for $n=0,1,2,\ldots,n_{max}$.
Table 1: Mass spectra of charmonium $M_{c\bar{c}}$ in GeV for $m_{c}=1.27$ GeV; $V_{1}=7.404662$ GeV; $V_{2}=-20.779639$ GeV; $V_{3}=14.959636$ GeV; $\alpha=0.223612$ GeV; $r_{e}=3.830931$ GeV-1. The percentage deviation from experimental values is presented below each calculated value.
@@
Table 2: Mass spectra of bottomonium $M_{b\bar{b}}$ in GeV for $m_{b}=4.18$ GeV; $V_{1}=4.883229$ GeV; $V_{2}=-12.928267$ GeV; $V_{3}=10.108982$ GeV; $\alpha=0.412272$ GeV; $r_{e}=1.836733$ GeV-1. The percentage deviation from experimental values is presented below each calculated value.
State Particle Present Work Theory I Theory II Theory III Experiment Non-Rel. Ref. [61] Ref. [62] Ref. [63] Ref. [75] 1S $\Upsilon$(1S) 9.460302 9.460 9.463 9.460 9.46030 0.00% 0.00% 0.03% 0.00% 1P $h_{b}$(1P) 9.594182 9.704 9.821 9.942 9.8993 -3.08% -1.97% -0.79% 0.43% 1D $\Upsilon_{2}$(1D) 9.827113 10.010 10.074 10.140 10.1637 -3.31% -1.51% -0.88% -0.23% 1F - 10.130188 10.268 10.288 - - 2S $\Upsilon$(2S) 10.023261 10.028 9.979 10.023 10.02326 0.00% 0.05% -0.44% 0.00% 2P $h_{b}$(2P) 10.151214 10.160 10.220 10.150 10.2598 -1.06% -0.97% -0.39% -1.07% 2D $\Upsilon_{2}$(2D) 10.380036 10.332 10.424 - - 2F - 10.688957 - 10.607 - - 3S $\Upsilon$(3S) 10.355200 10.343 10.359 10.178 10.3552 0.00% -0.12% 0.04% -1.71% 3P $h_{b}$(3P) 10.468567 - 10.556 - - 3D $\Upsilon_{2}$(3D) 10.681781 - 10.733 - - 3F - 10.986923 - 10.897 - - 4S $\Upsilon$(4S) 10.405400 10.536 10.683 10.242 10.5794 -1.64% -0.41% 0.98% -3.18% 4P $h_{b}$(4P) 10.497089 - 10.855 - - 4D $\Upsilon_{2}$(4D) 10.686643 - 11.015 - -
Briefly, in the spin symmetry, the GTHP’s inclusion of Woods–Saxon and Morse potentials means it can be used in nuclear bound-state problems (like nucleon–nucleus potential models) or molecular/atomic systems that employ Morse-type interactions. However, an equally important spin symmetry, pseudospin symmetry, also plays a crucial role in relativistic quantum systems [18, 19]. Pseudospin symmetry arises when the sum of the vector and scalar potentials satisfies $V(r)+S(r)=\text{constant}$[18, 19]. This symmetry is closely related to the near-degeneracy of nuclear energy levels and offers a complementary perspective to the spin symmetry analysis [25, 26]. The investigation of pseudospin symmetry is vital for understanding nuclear shell structures, meson spectroscopy, and the underlying physics governing quark-antiquark pairs [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. Therefore, we extend our analysis to pseudospin symmetry within the GTHP framework, examining its manifestation across different potential profiles and its impact on the system’s energy spectrum in the Section S2 of Ref. [79]. This transition from spin symmetry to pseudospin symmetry reinforces the unified nature of these symmetries and their relevance to nuclear and particle physics.